### v1

@@ 4.3 CNN-RetNet
In this section, we use CNN-RetNet to forecast PV power generation inspired by [38].
@@ 4.3.1 CNN
To extract the features of PV power generation, we add three convolutional layers before using RetNet. The first convolutional layer is used for the input data mapping. The second convolutional layer is used to extract features. The third convolutional layer is used to extract higher-order features.
@@ 4.3.2 Retention
where $\boldsymbol{B}\in\mathbb{R}^{d\times d}$ denotes a diagonal matrix, which is defined as $\boldsymbol{B}=\boldsymbol{\Lambda}(\gamma e^{i\theta})\boldsymbol{\Lambda}^{-1}$. With the equation (18), equation (19) and the definition of $\boldsymbol{\Lambda}$, we can rewrite the expression of $\boldsymbol{o}_{n}$ as follows:
@@
where $\boldsymbol{q}_{n}(\gamma e^{i\theta})^{n}$, $\boldsymbol{k}_{s}(\gamma e^{i\theta})^{-s}$ are relative position embeddings proposed by Transformer [39]. Moreover, when $\gamma$ is a scalar, equation (20) can be further simplified as follows:
@@
where $\odot$ denotes that the two matrices are multiplied element by element. $\Theta=\{\theta_{n}\}^{l_{x}}$ and $\overline{\Theta}$ denotes the complex conjugate of $\Theta$.
@@
We divide the input into chunks. We employ the parallel representation specified in equation (23) for computational processes within each chunk. Conversely, cross-chunk information is managed using the recurrent representation detailed in equation (24). $b$ denotes the chunk length. We calculate the retention output of the i-th chunk via:
@@
where the definition of $\boldsymbol{R}_{i}$ is similar to the state $\boldsymbol{s}_{n}$. These three forms of retention correspond to a, b, and c in figure 3 respectively.
@@ 4.3.3 Gated multi-scale retention
$h$ retention heads are employed in each layer, and $d_{h}$ satisfies $h=d/d_{h}$, where $d_{h}$ denotes head dimension. Each heads operates on distinct paramater matrices $\boldsymbol{W_{Q}},\boldsymbol{W_{K}},\boldsymbol{W_{V}}\in\mathbb{R}^{d\times d}$ and $\gamma$. Then to the input $\boldsymbol{X}$, the output of MSR layer is defined as follows:
@@ 4.3.4 Retention network
where gelu is the activation function.
@@
where $l\in[0,L]$. The output denotes $\boldsymbol{X}^{L}$. The process of RetNet is as follows:
@@ 4.4 Optimization
NSGA-II is used in multi-objective optimization. The main advantage is finding a balanced set of solutions. NSGA-II is applied in engineering, economics, logistics, and other fields. In this study, we adopt the NSGA-II algorithm.
@@ 5 Experiments
Figure 4: PV power generation forecasting using different models at the substation in January 2015
@@ 5.1 Experimental setup
5.1.1 Dataset
The dataset is multi-sample, multivariate and exhibiting temporal variation. The dataset originates from a PV power station in a specific region of Jiangsu Province, China, in 2015. We choose the dataset in January for the experiments because the PV data in Jiangsu is greatly affected by fog in January. The dataset involves 8 variables. Table 4 shows information about the variables.
5.1.2 Data pre-processing
We apply min-max normalization to the different scales of each variable. The model generates outputs from normalized input data, and outputs are inversely normalized for evaluation.
We choose data spanning 15 days to evaluate the model, maintaining a 3:1:1 ratio between the training, validating, and testing sets.
@@ 5.1.3 Implementation details
CCRetNet is used for 4-hour short-term forecasting. We set $m=5$, $\tau=2$ and $\beta=0.8$ in the calculation of TEWPP. We employ the silhouette coefficient to determine the optimal number of clusters. The hyperparameters are adjusted by employing the Optuna optimal framework.
Table 4: Variables and their units
Variable Name Unit
direct Normal Irradiance $W/m^{2}$
Global Horizontal Irradiance $W/m^{2}$
Diffuse Horizontal Irradiance $W/m^{2}$
Temperature of Component C
Ambient Temperature C
Atmospheric Pressure hPa
Relative Humidity %
PV Power Generation MW
@@ 5.1.4 Evaluation metrics
The objective of PV forecasting is to improve accuracy while reducing computational cost. Consequently, we use three evaluation metrics for accuracy and two for computational cost.
There are three evaluation metrics for accuracy: the Root Mean Squared Error (RMSE), the Mean Absolute Error (MAE), and the Coefficient of determination $R^{2}$. Each of them offers a unique perspective. RMSE focuses on larger values, MAE focuses on median error, and Coefficient of determination $R^{2}$ focuses on the proportion of variance in the dependent variable that is forecasting from the independent variables. The calculations of metrics are as follows:
@@
There are two evaluation metrics for computational cost: Floating Point Operations (FLOPs) and the number of parameters. FLOPs is used to measure the number of floating-point operations required by the model during the inference process. The number of parameters is used to evaluate the size and storage requirements of the model.
@@ 5.1.5 Computer environment
The experiments are conducted on two distinct setups. One includes a GeForce RTX 4090 GPU and a 13th Gen Intel(R) Core(TM) i9-13900KF processor on Ubuntu 22.04.3 LTS. Another includes a GeForce RTX 4060Ti GPU and a 13th Gen Intel(R) Core(TM) i5-13400F processor on Windows 11. Both configurations employ Python 3.10 and the Pytorch deep learning framework.
5.2 Results and analysis
We compare the proposed model with eight traditional models, including CCRetNet, CNN-Retnet, Retnet, CNN-LSTM, CNN-GRU, CNN-RNN, long short term memory (LSTM), gate recurrent unit (GRU), RNN, multilayer perceptron (MLP) and autoregressive integrated moving average model (ARIMA). We utilize the Optuna library for hyperparameters. We choose three methods for comparison: NSGA-II, random search (RS), and tree-structured parzen estimator (TPE). Model performance is evaluated based on accuracy and computational cost.
Fig 4 shows the forecasting results, and Fig 5 displays the numerical analysis. From the two subgraphs in Fig 4, CCRetNet provides the best forecasts when the real data fluctuates significantly. Fig 5 displays that CCRetNet reaches the lowest error values for RMSE and MAE, 2.71 MW and 2.55 MW, respectively, and the $R^{2}$ value is 0.97, higher than others. This illustrates the superiority of CCRetNet. ARIMA performs poorly in forecasting when the real data fluctuates significantly. Notably, the $R^{2}$ value of ARIMA is negative, indicating that ARIMA is unsuitable for PV forecasting on this dataset during the foggy winter. In the ablation experiments, the performance of CCRetNet is superior to CNN-RetNet due to clustering from Fig 4 and Fig 5. The performance of CNN-RetNet is better than that of RetNet due to CNN from Fig 4 and Fig 5. The performance of CNN is also supported by comparing other CNN-based deep learning models and general deep learning models.
Fig 6 shows the performance of the different optimization by three evaluation metrics for accuracy. The NSGA-II method outperforms the others, achieving the lowest RMSE and MAE and providing the most accurate forecasts. The RS method also shows commendable performance with a robust $R^{2}$ value, suggesting a reliable predictive capability. The TPE method appears less effective, as it exhibits the highest RMSE and MAE of others.
Figure 7 illustrates the computational cost associated with the three models. Among them, Retnet incurs the largest computational costs. Despite being slightly more computationally demanding than CNN-RetNet, CCRetNet significantly improves forecasting accuracy and trends during the foggy winter due to hierarchical clustering.
@@ 5.2 Results and analysis
Figure 5: The error of PV power generation forecasting using different models at the substation in January 2015.
Figure 6: The error of PV power generation forecasting using different optimization at the substation in January 2015
@@
Figure 7: The computational cost of PV power generation forecasting using different models at the substation in January 2015.
@@ 6 Conclusion
This study focuses on qualifying the uncertainty in a series of PV power generation during the foggy winter. The effectiveness of the proposed model is verified at a power station in a specific region of Jiangsu, China. We create TEWPP to quantify the uncertainty during the foggy winter. Additionally, we use a novel median linkage method for hierarchical clustering to reduce complexity. The optimal number of clusters $k$ is determined using the silhouette coefficient. Multiple CCRetNet models are trained and tested for various categories and hyperparameters optimized by the NSGA-II algorithm. The results indicate that CCRetNet outperforms other comparative models regarding RMSE, MAE, and $R^{2}$ for 4-hour short-term forecasting.
The study focuses on PV forecasting during the foggy winter without considering other weather scenarios. Therefore, our future research aims to create a more efficient and adaptable framework by focusing on year-round weather conditions to better forecast PV power generation across seasons.

### v2

@@ 4.3 CNN-RetNet
In this section, we use CNN-RetNet to make forecasts inspired by [32].
@@ 4.3.1 CNN
To extract features from PV power generation data, we add three convolutional layers before applying RetNet. The first convolutional layer maps the input data. The second layer extracts initial features, while the third layer captures higher-order features.
@@ 4.3.2 Retention
where $\boldsymbol{B}\in\mathbb{R}^{d\times d}$ denotes a diagonal matrix, defined as $\boldsymbol{B}=\boldsymbol{\Lambda}(\gamma e^{i\theta})\boldsymbol{\Lambda}^{-1}$. With the equation (18), equation (19) and the definition of $\boldsymbol{\Lambda}$, we can rewrite the expression of $\boldsymbol{o}_{n}$ as follows:
@@
where $\boldsymbol{q}_{n}(\gamma e^{i\theta})^{n}$, $\boldsymbol{k}_{s}(\gamma e^{i\theta})^{-s}$ are relative position embeddings proposed by Transformer [33]. Additionally, when $\gamma$ is a scalar, equation (20) can be further simplified as follows:
@@
where $\odot$ denotes the two matrices are multiplied element by element. $\Theta=\{\theta_{n}\}^{l_{x}}$ and $\overline{\Theta}$ denotes the complex conjugate of $\Theta$.
@@
We divide the input into chunks and apply the parallel representation specified in equation (23) for computations within each chunk. In contrast, cross-chunk information is handled using the recurrent representation described in equation (24). Here, $b$ represents the chunk length. The retention output for the $i$-th chunk is calculated as follows:
@@
where the definition of $\boldsymbol{R}_{i}$ is similar to the state $\boldsymbol{s}_{n}$. The three forms of retention correspond to a, b, and c in figure 3 respectively.
@@ 4.3.3 Gated multi-scale retention
$h$ retention heads are employed in each layer, and $d_{h}$ satisfies $h=d/d_{h}$, where $d_{h}$ denotes head dimension. Each heads operates on distinct paramater matrices $\boldsymbol{W_{Q}},\boldsymbol{W_{K}},\boldsymbol{W_{V}}\in\mathbb{R}^{d\times d}$ and $\gamma$. The output of the MSR layer for the input $\boldsymbol{X}$ is defined as follows:
@@ 4.3.4 Retention network
where gelu denotes the activation function.
@@
where $l\in[0,L]$. The output denotes $\boldsymbol{X}^{L}$. The process of RetNet is shown in algorithm 2.
@@ 4.4 Optimization
NSGA-II is used in multi-objective optimization and is known for finding a balanced set of solutions. It is commonly applied in engineering, economics, logistics, and other fields. In this study, we utilize the NSGA-II algorithm.
@@ 5.1 Experimental setup
5.1.1 Datasets
Both datasets are multi-sample, multivariate, and exhibit temporal variation. One dataset, which contains eight variables, originates from a 2015 PV power station in a specific region of Jiangsu Province, China. The other dataset originates from Beijing, China, collected in 2019 and contains nine variables. Table 4 details the variables in the 2015 Jiangsu dataset, while Table 5 outlines the variables in the 2019 Beijing dataset.
@@ 5.1.2 Implementation details
We apply min-max normalization to each variable at different scales. The model generates outputs from normalized input data, and outputs are inversely normalized for evaluation.
In calculating TEWPP, we set $m=5$, $\tau=2$, and $\beta=0.8$. We employ the silhouette coefficient to determine the optimal number of clusters. The detailed introduction of the silhouette coefficient has already been explained in Section 4.2.2. The hyperparameters are adjusted using the Optuna optimal framework.
Table 4: Statistical table of PV power generation and meteorological variables in a specific region of Jiangsu Province, 2015
Variable Name Unit Max Value Min Value Mean Value
Direct Normal Irradiance W/m2 1051.68 0.00 210.40
Global Horizontal Irradiance W/m2 962.29 0.00 192.52
Diffuse Horizontal Irradiance W/m2 94.65 0.00 18.94
Temperature of Component C 53.00 -6.30 16.65
Ambient Temperature C 36.50 -4.70 15.44
Atmospheric Pressure hPa 1040.40 994.00 1017.52
Relative Humidity % 100.00 15.00 76.82
PV Power Generation MW 115.775 -0.44 17.68
Table 5: Statistical table of PV power generation and meteorological variables in Beijing, 2019
Variable Name Unit Max Value Min Value Mean Value
Air Temperature C 40.973 -16.618 12.922
Precipitation mm/hour 7.493 0.000 0.056
Snowfall mm/hour 3.242 0.000 0.002
Snow Mass kg/m2 7.001 0.000 0.148
Air Density kg/m3 1.378 1.093 1.212
Ground-level Solar Irradiance W/m2 1012.688 0.000 191.877
Top of Atmosphere Solar Irradiance W/m2 1259.022 0.000 328.785
Cloud Cover Fraction [0, 1] 0.999 0.000 0.351
PV Power Generation MW 0.875 0.000 0.177
@@ 5.1.3 Evaluation metrics
The objective of PV forecasting is to improve accuracy while reducing computational costs. Consequently, we use three evaluation metrics for accuracy and two for computational costs.
There are three evaluation metrics for accuracy: the Root Mean Squared Error (RMSE), the Mean Absolute Error (MAE), and the coefficient of determination $R^{2}$. Each of them offers a unique perspective. RMSE focuses on larger values, MAE focuses on median error, and Coefficient of determination $R^{2}$ focuses on the proportion of variance in the dependent variable forecasting from the independent variables. The calculations of metrics are as follows:
@@
There are two evaluation metrics for computational costs: floating-point operations (FLOPs) and the number of parameters. FLOPs measure the number of floating-point operations required by the model during the inference process, while the number of parameters evaluates the model’s size and storage requirements.
@@ 5.1.4 Computer environment
The experiments are conducted on two distinct setups. One includes a GeForce RTX 4090 GPU and a 13th Gen Intel(R) Core(TM) i9-13900KF processor on Ubuntu 22.04.3 LTS. Another includes a GeForce RTX 4060Ti GPU and a 13th Gen Intel(R) Core(TM) i5-13400F processor on Windows 11. Both configurations use Python 3.10 and the Pytorch deep learning framework.
5.2 Analysis of air quality
In this section, we analyze the air quality of a specific region in Jiangsu Province in 2015 and the air quality in Beijing in 2019. The air quality datasets for both locations can be found at https://www.aqistudy.cn/historydata/.
Fig. LABEL:AQIyan and LABEL:AQIbeijing show the monthly air quality days for a region in Jiangsu and Beijing, respectively. It can be observed that there are more days with severe pollution in January and December, while the air quality is better in the summer. Fig. LABEL:PM2.5yan and LABEL:PM2.5beijing present the monthly variation of PM2.5 concentrations in the same regions. PM2.5 concentrations fluctuate significantly in winter while they remain relatively stable in summer.
From the four figures, we can conclude that air quality in both datasets exhibits significant seasonal variations, with more severe pollution levels typically occurring during the winter months, particularly in January. The data shows that PM2.5 concentrations in January are highly variable, with a wide range of values and numerous outliers, indicating unstable and often high pollution levels.
Based on the above analysis, we selected the January data from the two datasets for single-step forecasting experiments. For multi-step forecasting experiments, we used the Beijing data from January to March, where the first 12 points were used to predict the subsequent 4 points.
@@ 5.2 Analysis of air quality
Figure 7: Comparison of PV power generation forecasting on different models in a specific region of Jiangsu Province, January 2015
5.3 Experimental results and analysis
We conduct experiments using two datasets, with the results discussed in Section 5.3.1, Section 5.3.2, and Section 5.3.3.
We compare the proposed model with eight traditional models, including CCRetNet, CNN-Retnet, Retnet, CNN-LSTM, CNN-GRU, CNN-RNN, long short term memory (LSTM), gate recurrent unit (GRU), recurrent neural network (RNN), autoregressive integrated moving average model (ARIMA) and Transformer. We utilize the Optuna library for hyperparameters and choose three methods for comparison: NSGA-II, random search (RS), and tree-structured parzen estimator (TPE). Model performance is evaluated based on accuracy and computational costs.
5.3.1 Analysis of single-step forecasting experiments in Jiangsu Province, 2015
In this section, we emphasize the experimental outcomes of multiple forecasting models applied to a specific region of Jiangsu Province dataset.
Figure 9: The computational costs on different models in a specific region of Jiangsu Province, January 2015
We compare the forecasting performance of various models, highlighting the superiority of CCRetNet in handling significantly fluctuating real data. Fig. 7 presents the forecasting results, while Fig. LABEL:error provides numerical analysis. CCRetNet performs best when real data fluctuates considerably, with the lowest RMSE and MAE values of 2.71 MW and 2.55 MW and $R^{2}$ value of 0.97, higher than the other models, demonstrating its superiority. In contrast, ARIMA performs poorly when real data fluctuates significantly, with a negative $R^{2}$ value, indicating that it is unsuitable for PV forecasting on this dataset, especially during hazy weather.
The ablation experiments further verify the superiority of CCRetNet. Fig. 7 and Fig. LABEL:error show that CCRetNet outperforms CNN-RetNet, primarily due to the clustering method based on TEWPP, which effectively qualifies uncertainty during hazy weather and classifies them within the clustering module. Fig. 7 and Fig. LABEL:error also indicate that CNN-RetNet performs better than RetNet due to the introduction of CNN. The effectiveness of CNN is further supported by comparisons with other CNN-based deep learning models and general deep learning models.
Fig. LABEL:optimize shows the accuracy of CCRetNet under different optimization methods across three evaluation metrics. The model optimized by NSGA-II outperforms the others. The model optimized by RS also shows commendable performance with a robust $R^{2}$ value, suggesting a reliable predictive capability. The model optimized by TPE appears less effective, as it exhibits the highest RMSE and MAE of others.
Fig. 9 illustrates the computational costs of the three models. RetNet incurs the highest computational costs. Although CCRetNet’s computational costs are slightly higher than that of CNN-RetNet, the introduction of hierarchical clustering significantly improves CCRetNet’s forecasting accuracy and trend performance.
5.3.2 Analysis of single-step forecasting experiments in Beijing, 2019
@@
Figure 10: Comparison of PV power generation forecasting on different models in Beijing, January 2019
This section presents the performance analysis of various forecasting models on the Beijing dataset, highlighting both trend alignment and numerical accuracy.
Fig. 10 compares the actual values and the forecasts of traditional models on the Beijing dataset for January 2019. Overall, the $R^{2}$ values indicate that traditional models perform better on the Beijing dataset than on the Jiangsu dataset. CCRetNet exhibits minimal deviation from actual results during stable periods, although its performance during peak values is comparable to other traditional models as shown in Fig. 10.
Moreover, Fig. LABEL:error1_beijingsingle reveals that CCRetNet achieves the smallest deviation from actual results, with the ablation experiment further confirming its superiority over CNN-RetNet, particularly due to the effectiveness of TEWPP-based clustering in qualifying uncertainty.
Fig. LABEL:optimize_beijingsingle compares the performance of CCRetNet under different optimization methods. The model optimized by NSGA-II achieves the best forecasting accuracy, evidenced by the lowest RMSE and MAE. The models optimized by TPE and RS both outperform the NSGA-II optimized model. However, they show similar forecasting accuracy on the Beijing dataset, which differs from the results on the Jiangsu dataset.
Figure 11: Comparison of PV power generation forecasting on different models in Beijing, March 2019
5.3.3 Analysis of multi-step forecasting experiments in Beijing, 2019
In this section, we present the experimental results of several multi-step forecasting models on the Beijing dataset.
In multi-step forecasting experiments, most models perform worse than in single-step forecasting, as shown in Fig. 11, Fig. LABEL:error1_beijingmulti, and Fig. LABEL:optimize_beijingmulti. Although the MSE and MAE of CNN-BiLSTM and Transformer are higher in multi-step forecasting, the $R^{2}$ is closer to 1 compared to single-step forecasting.
In the ablation experiments, CCRetNet exhibits smaller forecasting fluctuations compared to other models when PV power generation is stable, similar to the results of single-step forecasting. The ablation experiments also demonstrate the superiority of the clustering method based on TEWPP and CNN method.
The CCRetNet model optimized by NSGA-II performs the best in multi-step forecasting, comparing optimization methods. As shown in Fig. LABEL:optimize_beijingmulti, the model optimized by TPE shows better forecasting accuracy than the model optimized by RS. However, both are outperformed by the model optimized by NSGA-II, a result that contrasts with the previous two experiments.
@@ 6 Conclusion
This study focuses on qualifying uncertainty in a series of PV power generation during hazy weather. The effectiveness of the proposed model is validated with two datasets. We create TEWPP to qualify uncertainty during hazy weather. Additionally, we use a novel median linkage method for hierarchical clustering to reduce computational costs. The optimal number of clusters $k$ is determined using the silhouette coefficient. Multiple CCRetNet models are trained and tested for various categories and hyperparameters optimized by the NSGA-II algorithm. The results indicate that CCRetNet outperforms other comparative models regarding RMSE, MAE, and $R^{2}$ for short-term forecasting.
The study focuses on PV forecasting during hazy weather without considering other weather scenarios. Therefore, our future research aims to create a more efficient and adaptable framework by focusing on year-round weather conditions to make forecasts better.